# The special unitary groups $SU(2n)$ as framed manifolds

### v3

@@ Abstract.
Let $[SU(2n),\mathscr{L}]$ denote the bordism class of $SU(2n)$ $(n\geq 2)$ equipped with the left invariant framing $\mathscr{L}$. Then it is well known that $e_{\mathbb{C}}([SU(2n),\mathscr{L}])=0$ in $\mathbb{O}/\mathbb{Z}$ where $e_{\mathbb{C}}$ denotes the complex Adams $e$-invariant. In this note we show that replacing $\mathscr{L}$ by the twisted framing by a specific map it can be transformed into a generator of $\mathrm{Im}\,e_{\mathbb{C}}$. In addition to that we also show that the same procedure affords an analogous result for a quotient of $SU(2n+1)$ by a circle subgroup which inherits a canonical framing from $SU(2n+1)$ in the usual way.
@@ 1. Introduction
where $e_{\mathbb{C}}:\pi_{4\mathit{l}-1}^{S}\to\mathbb{Q}/\mathbb{Z}$ denotes the complex Adams $e$-invariant. In view of this result we wish to raise the problem of the existence of a map $f:G\to GL(s,\mathbb{R})$ such that $e_{\mathbb{C}}([G,\mathscr{L}^{f}])$ is nonzero and if possible, is equal to $(-1)^{\ell-1}B_{\ell}/2\ell$  [2] where $\mathscr{L}^{f}$ denotes the framing obtained by twisting $\mathscr{L}$ by $f$ and $B_{\mathit{l}}$ the $\mathit{l}$th Bernoulli number.
In this note we consider the case $G=SU(2n)$ $(n\geq 2)$ where $\dim G=4n^{2}-1$. The main result is the following theorem.
@@
The proof is based on Proposition 2.1 of  [5]. But also in the case $G=SU(2n+1)$ it can be seen that the procedure used there is adaptable to its quotient by a circle subgroup $C$, considered as a framed manifold equipped with the induced framing $\mathscr{L}_{C}$. Now this framing can be obtained by applying the fact that $\mathscr{L}$ is a $C$-equivariant framing  [4] to the equation ()($\ast$)( ∗ ) of  [5, p. 16]. Here let $S^{1}=\{z\in\mathbb{C}||z|=1\}$ and put
@@
where $\mathrm{diag}(a_{1},\ldots,a_{2n+1})$ denotes the diagonal matrix with $a_{1},\ldots,a_{2n+1}$ on the diagonal. Then in fact we have
@@ 2. Tensor product decomposition of $E$
for the elements of $SU(2)$. Then $R(rz,v)d(z)=R(r,zv)$ and so letting it correspond to the element $(1-2r^{2},2rzv)\in S^{2}$ we have a homeomorphism between $SU(2)/S$ and $S^{2}$. Thinking of $p$ as a principal $S$-bundle over $S^{2}$ via this homeomorphism we have $(S(L),\pi,S^{2})\cong(SU(2),p,S^{2})$ where $L$ is used instead of $E$. From the formula of  [5, p. 40] we have $e([S(L),\Phi_{L}])=B_{1}/4$ where $\Phi_{L}$ denotes the trivialization of the stable tangent bundle of $S(L)$. In this case $\Phi_{L}$ coincides with $\mathscr{L}$ via the above isomorphism and so we have $e([SU(2),\mathscr{L}])=1/24$.
@@
$x=(x_{0},\cdots,x_{m-2})\to p\biggl{(}\textstyle\prod_{j=0}^{m-2}R_{j}(r_{i;j},w_{i;j})\biggr{)}$
where $x_{j}=(x_{1;j},\cdots,x_{m_{j};j})\in(S_{j}^{2})^{m_{j}}$ with $x_{i;j}=(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j}$. If we put $R(r_{i;j}z_{i},v_{i;j})=\prod_{j=0}^{m-2}R_{j}(r_{i;j}z_{i},v_{i;j})$, then as in the case of $P^{\{i\}}_{j}$, using (5), (5b), (6) we see that
@@
forms the total space of a principal $S$-bundle endowed with the projection map $q:P\to(S^{2})^{(m^{2}-m)/2}$ such that $\phi\circ q=p|P$. If we write $L^{(m^{2}-m)/2}=L_{0}^{m_{0}}\boxtimes\cdots\boxtimes L_{m-2}^{m_{m-2}}$ where $L_{j}^{m_{j}}=L_{1;j}\boxtimes\cdots\boxtimes L_{m_{j};j}$, then $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong(P,q,(S^{2})^{(m^{2}-m)/2})$ for the reason similar to (7). Therfore we have
(8) $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong\phi^{*}(G,p,G/S)$
where $\phi^{*}(-)$ denotes the induced bundle.
@@
$R_{0}(r_{i;0},w_{i;0})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{0}(r^{\prime}_{i;0},w^{\prime}_{i;0})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
@@
Substituting the result of (3) into these equations we have $r_{1;0}\cdots r_{m-1;0}=r^{\prime}_{1;0}\cdots r^{\prime}_{m-1;0}$, $r_{s+1;0}\cdots r_{m-1;0}\bar{w}_{s;0}=r^{\prime}_{s+1;0}\cdots r^{\prime}_{m-1;0}\bar{w}^{\prime}_{s;0}$ $(1\leq s\leq m-2)$ and $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ where $r_{s;0}^{2}+|w_{s;0}|^{2}=1$ and $r_{s;0}\geq 0$. For convenience we assume that $r_{i;0}>0$ and moreover proceed by induction on $i$ in reverse order. First from $\underline{a}_{m-1;0}=\underline{a}^{\prime}_{m-1;0}$ we have $w_{m-1;0}=w^{\prime}_{m-1;0}$ and so it follows that $r_{m-1;0}=r^{\prime}_{m-1;0}$. Next from the second equation $\underline{a}_{m-2;0}=\underline{a}^{\prime}_{m-2;0}$ we see that $r_{m-1;0}w_{m-2;0}=r^{\prime}_{m-1;0}w^{\prime}_{m-2;0}$. But since $r_{m-1;0}=r^{\prime}_{m-1;0}>0$ we have $w_{m-2;0}=w^{\prime}_{m-2;0}$ and so similarly it follows that $r_{m-2;0}=r^{\prime}_{m-2;0}$. Repeating this procedure we obtain subsequently $w_{m-3;0}=w^{\prime}_{m-3;0}$, $r_{m-3;0}=r^{\prime}_{m-3;0}$, $w_{m-4;0}=w^{\prime}_{m-4;0}$, $r_{m-4;0}=r^{\prime}_{m-4;0}$$\cdots$, $r_{2;0}=r^{\prime}_{2;0}$ and finally we obtain $w_{1;0}=w^{\prime}_{1;0}$, $r_{1;0}=r^{\prime}_{1;0}$. At this point we conclude that $x_{0}=x^{\prime}_{0}$ and so the assumption above can be rewritten as
@@
with the same condition as above. Repeating this step we obtain $x_{2}=x^{\prime}_{2}$, $\ldots$, $x_{m-2}=x^{\prime}_{m-2}$ via inductive reduction, which implies $x=x^{\prime}$. This proves the lemma.
In the above, if $r_{m_{j};j}=0$, then we think of it as not being zero by exchanging the $m_{j}$th and other proper rows and columns depending on the case. In the case that there are no such row and column, this approach leads us to obtain directly $r_{i;j}=r^{\prime}_{i;j}=0$ $(1\leq i\leq m_{j})$ and so $x_{j}=x^{\prime}_{j}$. ∎
@@ 3. Proofs of Theorem and Proposition
for $z_{2k-1},z_{2k},z\in S^{1}$. Let $Q_{k}$ be the subspace of $G$ consisting of $D(z_{2k-1},zz_{2k})d(\bar{z})$. Then it forms the total space of a principal $S$-bundle over $T_{k}=S^{1}\times S^{1}$ along with the projective map $p_{k}:Q_{k}\to T_{k}$ given by $D(z_{2k-1},zz_{2k})d(\bar{z})\to(z_{2k-1},zz_{2k})$ where $T_{k}$ is considered as an subspace of $G/S$ under $\psi_{k}:(z_{2k-1},zz_{2k})\to p(D(z_{2k-1},zz_{2k}))$.
@@
$\mu(e^{\eta i},ze^{\theta i})=\left\{\begin{array}[]{ll}(\cos(\eta/2),ze^{\theta i}\sin(\eta/2))_{R}&\ (0\leq\eta\leq\pi)\\ (-\cos(\eta/2),z\sin(\eta/2))_{R}&\ (\pi\leq\eta<2\pi).\end{array}\right.$
Then taking into account the fact that a principal circle bundle over $S^{1}$ is trivial we see that the classifying map of $p_{k}$ factors through $S^{2}$ and therefore we have
Lemma 2.
$(Q_{k},p_{k},T_{k})\cong\mu^{*}(SU(2),p,S^{2})$ and $\mu^{*}:H^{2}(S^{2},\mathbb{Z})\stackrel{{\scriptstyle\cong}}{{\to}}H^{2}(T_{k},\mathbb{Z})$
$(1\leq k\leq m/2-1)$.
@@
$\tilde{\mu}(D(e^{\eta i},ze^{\theta i})d(\bar{z}))=\left\{\begin{array}[]{ll}R(\bar{z}\cos(\eta/2),e^{\theta i}\sin(\eta/2))&\ (0\leq\eta\leq\pi)\\ R(-\bar{z}\cos(\eta/2),\sin(\eta/2))&\ (\pi\leq\eta<2\pi),\end{array}\right.$
@@
Then as in the case of $P$
@@
forms the total space of a principal $S$-bundle endowed with the projection map $\tilde{q}:\tilde{P}\to(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ such that $\psi\circ\tilde{q}=p|P$. Let $\tilde{L}_{k}$ be the complex line bundle over $T_{k}$ associated to $p_{k}$. Write $\tilde{L}^{n-1}=\tilde{L}_{1}\boxtimes\cdots\boxtimes\tilde{L}_{n-1}$ and put $E=L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$. Then $(S(E),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong(\tilde{P},q,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})$ and therefore we have
(9) $(S(E),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong\psi^{*}(G,p,G/S).$
@@
$\psi:(S^{2})^{(m^{2}-m)/2}\times T^{n-1}\to G/S$ is injective.
@@
In terms of the notation of the proof of Lemma 1 we prove that if we suppose that $\psi(x,y)=\psi(x^{\prime},y^{\prime})$, then it follows that $(r_{i;j},w_{i;j})_{R}=(r^{\prime}_{i;j},w^{\prime}_{i;j})_{R}$ $(0\leq j\leq m-2)$ and $(z_{2k-1},\zeta_{2k})=(z^{\prime}_{2k-1},\zeta^{\prime}_{2k})$ $(1\leq k\leq n-1)$, the latter of which is the part added to the proof of injectivity of $\phi$. The proof can be proceeded along the same lines as in the proof of the injectivity of $\phi$ above. But in fact there are the following changes in the use of the elements $\bar{w}_{m_{j};j}$ each of which is a starting point of the proof for $\phi_{j}$: $\bar{w}_{m-1;0}\to\bar{w}_{m-1;0}$ (i.e. no change), $\bar{w}_{m-2:1}\to z_{1}\bar{w}_{m-2:1}$, $\bar{w}_{m-3:2}\to(\bar{z}_{1}\zeta_{2})\bar{w}_{m-3:2}$ and in general $\bar{w}_{m-2k-2;2k+1}\to(z_{2k+1}\bar{\zeta}_{2k})\bar{w}_{m-2k-2;2k+1}$, $\bar{w}_{m-2k-3;2k+2}\to(\bar{z}_{2k+1}\zeta_{2k+2})\bar{w}_{m-2k-3;2k+2}$ for $k=1,\ldots,n-2$. Similarly substituting these elements into the equations $\underline{a}_{0;j}=\underline{a^{\prime}}_{0;j}$, $\underline{a}_{1;j}=\underline{a^{\prime}}_{1;j}$, $\ldots$, $\underline{a}_{m-j-1;j}=\underline{a^{\prime}}_{m-j-1;j}$ we can obtain $(x,y)=(x^{\prime},y^{\prime})$.
First from $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ we have $x_{0}=x^{\prime}_{0}$, which shows that $z_{1}\bar{w}_{m-2:1}=z^{\prime}_{1}\bar{w}^{\prime}_{m-2:1}$, so we have $z_{1}r_{m-2:1}=z^{\prime}_{1}r^{\prime}_{m-2:1}$. If $r_{m-2:1}>0$, then from the second equation it follows that $r_{m-2:1}=r^{\prime}_{m-2:1}$ and so $z_{1}=z^{\prime}_{1}$. If $r_{m-2:1}=0$, then since $x_{0}=x^{\prime}_{0}=(0,1)_{R}$ are satisfied we get $z_{1}=z^{\prime}_{1}$. However in this case in order to move the next step we need to use the approrch of exchanging rows and columns. In any case it can be concluded that $x_{1}=x^{\prime}_{1}$. $z_{1}=z^{\prime}_{1}$, Next from $(\bar{z}_{1}\zeta_{2})\bar{w}_{m-3:2}=(\bar{z}^{\prime}_{1}\zeta^{\prime}_{2})\bar{w}^{\prime}_{m-3:2}$ we have $\zeta_{2}\bar{w}_{m-3:2}=\zeta^{\prime}_{2}\bar{w}^{\prime}_{m-3:2}$ and so it follows that $x_{2}=x^{\prime}_{2}$, $\zeta_{2}=\zeta^{\prime}_{2}$. Therefore by $(z_{3}\bar{\zeta}_{2})\bar{w}_{m-4;3}=(z^{\prime}_{3}\bar{\zeta}^{\prime}_{2})\bar{w}^{\prime}_{m-4;3}$ we have $z_{3}\bar{w}_{m-4;3}=z^{\prime}_{3}\bar{w}^{\prime}_{m-4;3}$ and so in a similar way to the previous case we have $x_{3}=x^{\prime}_{3}$, $z_{3}=z^{\prime}_{3}$. Subsequently also by $(\bar{z}_{3}\zeta_{4})\bar{w}_{m-5;4}=(\bar{z}^{\prime}_{3}\zeta^{\prime}_{4})\bar{w}^{\prime}_{m-5;4}$ we have $\zeta_{4}\bar{w}_{m-5;4}=\zeta^{\prime}_{4}\bar{w}^{\prime}_{m-5;4}$ and so it follows that $x_{4}=x^{\prime}_{4}$, $\zeta_{4}=\zeta^{\prime}_{4}$. At this point we can conclude that $x_{1}=x^{\prime}_{1}$, $x_{2}=x^{\prime}_{2}$, $x_{3}=x^{\prime}_{3}$, $x_{4}=x^{\prime}_{4}$ and $y_{1}=y^{\prime}_{1}$, $y_{2}=y^{\prime}_{2}$. These tell us that through repetition of these procedures we can arrive at the desired result. ∎
@@
First we note that $\dim((S^{2})^{(m^{2}-m)/2}\times T^{n-1})=\dim G/S$. In general $SU(m)$ is embedded in $\mathbb{C}^{m(m-1)}$ as the subspace of $(m-1)$-tuples $(x_{1},\cdots,x_{m-1})$ with $x_{i}\in\mathbb{C}^{m}$, $|x_{i}|=1$ and $x_{i}\bot x_{j}$ if $i\neq j$ (cf.  [1] and  [6]). In view of this, if we think of $\psi$ as a map between manifolds of the same dimension $m^{2}-2$, then by taking into account its differentiability we can see from Lemma 3 that $\psi$ is a homeomorphism and so by construction it can also be taken to be a diffeomorphism.
Identify $G/S$ with $(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ via $\psi$ and write $E$ for $L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$ in (9). Then combining the equations of (7) and Lemma 2 we have from Proposition 2.1 of [5]

### v4

@@ Abstract.
Let $[SU(2n),\mathscr{L}]$ denote the bordism class of $SU(2n)$ $(n\geq 2)$ equipped with its left invariant framing $\mathscr{L}$. Then it is well known that $e_{\mathbb{C}}([SU(2n),\mathscr{L}])=0$ where $e_{\mathbb{C}}$ denotes the complex Adams $e$-invariant. In this note we show that replacing $\mathscr{L}$ by the framing obtained by twisting it by a specific map the zero value of $e_{\mathbb{C}}([SU(2n),\mathscr{L}])$ can be transformed into a generator of $\mathrm{Im}\,e_{\mathbb{C}}$ which is isomorphic to a cyclic group. In addition we show that the same procedure affords an analogous result for a quotient of $SU(2n+1)$ by a circle subgroup which inherits a canonical framing from $SU(2n+1)$ in the usual way.
@@ 1. Introduction
where $e_{\mathbb{C}}:\pi_{4\mathit{l}-1}^{S}\to\mathbb{Q}/\mathbb{Z}$ denotes the complex Adams $e$-invariant. In view of this result we wish to raise the problem of the existence of a map $f:G\to GL(s,\mathbb{R})$ such that $e_{\mathbb{C}}([G,\mathscr{L}^{f}])$ is nonzero and if possible, is equal to $(-1)^{l-1}B_{l}/2l$  [1] where $\mathscr{L}^{f}$ denotes the framing obtained by twisting $\mathscr{L}$ by $f$ ( [2], [5], [7]) and $B_{l}$ the $l$th Bernoulli number.
In this note we consider the case $G=SU(2n)$ with dimension $14n^{2}-$ 1 $(n\geq 2)$. The main result is the following theorem.
@@
The proof is based on Proposition 2.1 of  [6]. But also in the case $G=SU(2n+1)$ it can be seen that the procedure used there is adaptable to its quotient by a circle subgroup $C$, considered as a framed manifold equipped with the induced framing $\mathscr{L}_{C}$. Now this framing can be obtained by applying the fact that $\mathscr{L}$ is a $C$-equivariant framing  [5] to the equation ()($\ast$)( ∗ ) of  [6, p. 16]. Let $S^{1}=\{z\in\mathbb{C}||z|=1\}$ and put
@@
where $\mathrm{diag}(a_{1},\ldots,a_{2n+1})$ denotes the diagonal matrix with diagonal entries $a_{1},\ldots,a_{2n+1}$. Then we have
@@ 2. Tensor product decomposition of $E$
for the elements of $SU(2)$. Then $R(rz,v)d(z)=R(r,zv)$ and so letting it correspond to the element $(1-2r^{2},2rzv)\in S^{2}$ we have a homeomorphism between $SU(2)/S$ and $S^{2}$. Thinking of $p$ as a principal $S$-bundle over $S^{2}$ via this homeomorphism we have $(S(L),\pi,S^{2})\cong(SU(2),p,S^{2})$ where $L$ is used instead of $E$. From the formula of  [6, p. 40] we have $e([S(L),\Phi_{L}])=B_{1}/4$ where $\Phi_{L}$ denotes the trivialization of the stable tangent bundle of $S(L)$. In this case $\Phi_{L}$ coincides with $\mathscr{L}$ via the above isomorphism and so we have $e([SU(2),\mathscr{L}])=1/24$.
@@
$x=(x_{0},\ldots,x_{m-2})\to p\biggl{(}\textstyle\prod_{j=0}^{m-2}R_{j}(r_{i;j},w_{i;j})\biggr{)}$
where $x_{j}=(x_{1;j},\ldots,x_{m_{j};j})\in(S_{j}^{2})^{m_{j}}$ with $x_{i;j}=(r_{i;j},w_{i;j})_{R}\in S^{2}_{i;j}$. If we put $R(r_{i;j}z_{i},v_{i;j})=\prod_{j=0}^{m-2}R_{j}(r_{i;j}z_{i},v_{i;j})$, then as in the case of $P^{\{i\}}_{j}$, using (5), (5b), (6) we see that
@@
forms the total space of a principal $S$-bundle endowed with the projection map $q:P\to(S^{2})^{(m^{2}-m)/2}$ such that $\phi\circ q=p|P$. If we write $L^{(m^{2}-m)/2}=L_{0}^{m_{0}}\boxtimes\cdots\boxtimes L_{m-2}^{m_{m-2}}$ where $L_{j}^{m_{j}}=L_{1;j}\boxtimes\cdots\boxtimes L_{m_{j};j}$, then $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong(P,q,(S^{2})^{(m^{2}-m)/2})$ for the reason similar to (7). Therefore we have
(8) $(S(L^{(m^{2}-m)/2}),\pi,(S^{2})^{(m^{2}-m)/2})\cong\phi^{*}(S(E),\pi,G/S),$
where $\phi^{*}(-)$ denotes the induced bundle by $\phi$.
Let $(S^{2}_{i;j})^{\circ}\in S^{2}_{i;j}$ be the subspace consisting of $(r_{i;j},w_{i;j})_{R}$ with $r_{i;j}>0$, and put $((S^{2})^{m_{j}})^{\circ}=\prod_{i=1}^{m_{j}}(S^{2}_{i;j})^{\circ}$ and $((S^{2})^{(m^{2}-m)/2})^{\circ}=\prod_{j=0}^{m-2}((S^{2})^{m_{j}})^{\circ}$. Then
@@
($\ast$) $R_{0}(r_{i;0},w_{i;0})\cdots R_{m-2}(r_{i;m-2},w_{i;m-2})=R_{0}(r^{\prime}_{i;0},w^{\prime}_{i;0})\cdots R_{m-2}(r^{\prime}_{i;m-2},w^{\prime}_{i;m-2})$
@@
Substituting the result of (3) into these equations we have $r_{1;0}\cdots r_{m-1;0}=r^{\prime}_{1;0}\cdots r^{\prime}_{m-1;0}$, $r_{s+1;0}\cdots r_{m-1;0}\bar{w}_{s;0}=r^{\prime}_{s+1;0}\cdots r^{\prime}_{m-1;0}\bar{w}^{\prime}_{s;0}$ $(1\leq s\leq m-2)$ and $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ where $r_{s;0}^{2}+|w_{s;0}|^{2}=1$ and $r_{s;0}>0$ by the assumption. We now proceed by induction on $i$ in reverse order. First from $\underline{a}_{m-1;0}=\underline{a}^{\prime}_{m-1;0}$ we have $w_{m-1;0}=w^{\prime}_{m-1;0}$ and so it follows that $r_{m-1;0}=r^{\prime}_{m-1;0}$. Next from the second equation $\underline{a}_{m-2;0}=\underline{a}^{\prime}_{m-2;0}$ we see that $r_{m-1;0}w_{m-2;0}=r^{\prime}_{m-1;0}w^{\prime}_{m-2;0}$. But since $r_{m-1;0}=r^{\prime}_{m-1;0}>0$ we have $w_{m-2;0}=w^{\prime}_{m-2;0}$ and so similarly it follows that $r_{m-2;0}=r^{\prime}_{m-2;0}$. Repeating this procedure we obtain subsequently $w_{m-3;0}=w^{\prime}_{m-3;0}$, $r_{m-3;0}=r^{\prime}_{m-3;0}$, $w_{m-4;0}=w^{\prime}_{m-4;0}$, $r_{m-4;0}=r^{\prime}_{m-4;0}$$\cdots$, $r_{2;0}=r^{\prime}_{2;0}$ and finally we obtain $w_{1;0}=w^{\prime}_{1;0}$, $r_{1;0}=r^{\prime}_{1;0}$. At this point we conclude that $x_{0}=x^{\prime}_{0}$ and so the equation ()($\ast$)( ∗ ) above can be rewritten as
@@
with the same condition as above. Repeating this step we obtain $x_{2}=x^{\prime}_{2}$, $\ldots$, $x_{m-2}=x^{\prime}_{m-2}$ via inductive reduction, which implies $x=x^{\prime}$. This proves the lemma. ∎
@@ 3. Proofs of Theorem and Proposition
for $z_{2k-1},z_{2k},z\in S^{1}$. Let $Q_{k}$ be the subspace of $G$ consisting of $D(z_{2k-1},zz_{2k})d(\bar{z})$. Then it forms the total space of a principal $S$-bundle over $T_{k}=S^{1}\times S^{1}$ along with the projective map $p_{k}:Q_{k}\to T_{k}$ given by $D(z_{2k-1},zz_{2k})d(\bar{z})\to(z_{2k-1},zz_{2k})$ where $T_{k}$ is considered as a subspace of $G/S$ under $\psi_{k}:(z_{2k-1},zz_{2k})\to p(D(z_{2k-1},zz_{2k}))$.
@@
$\mu(e^{\eta i},ze^{\theta i})=\left\{\begin{array}[]{ll}(\cos(\eta/2),ze^{\theta i}\sin(\eta/2))_{R}&\ (0\leq\eta\leq\pi)\\ (-\cos(\eta/2),ze^{\theta ti}\sin(\eta/2))_{R}&\ (\pi\leq\eta<2\pi),\ \ \ t=2-\eta/\pi.\end{array}\right.$
Then taking into account the fact that a principal circle bundle over $S^{1}$ is trivial we see that the classifying map of $p_{k}$ factors through $S^{2}$ where the restriction of $p_{k}$ to $\{1\}\times S^{1}\subset T_{k}$ is viewed as being trivial. Therefore we have
Lemma 2 (cf. [6], §2, Example 3).
$(Q_{k},p_{k},T_{k})\cong\mu^{*}(SU(2),p,S^{2})$ and also $\mu$ induces an isomorphism $H^{2}(S^{2},\mathbb{Z})\cong H^{2}(T_{k},\mathbb{Z})$ $(1\leq k\leq m/2-1)$.
@@
$\tilde{\mu}(D(e^{\eta i},ze^{\theta i})d(\bar{z}))=\left\{\begin{array}[]{ll}R(\bar{z}\cos(\eta/2),e^{\theta i}\sin(\eta/2))&\ (0\leq\eta\leq\pi)\\ R(-\bar{z}\cos(\eta/2),e^{\theta ti}\sin(\eta/2))&\ (\pi\leq\eta<2\pi),\ \ \ t=2-\eta/\pi,\end{array}\right.$
@@
Here we see by definition that the product terms in the second-half of the right-hand side satisfy $d(\bar{z}^{\prime})D(z_{2k-1},z{z^{\prime}}^{2}z_{2k})d(\bar{z})=D(z_{2k-1},zz_{2k})d(\bar{z})d(\bar{z}^{\prime})$. Using these equations together with the ones of (5), (5b) and (6), as in the case of $P$ above, we see that
@@
forms the total space of a principal $S$-bundle endowed with the projection map $\tilde{q}:\tilde{P}\to(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$ such that $\psi\circ\tilde{q}=p|\tilde{P}$. Let $\tilde{L}_{k}$ be the complex line bundle over $T_{k}$ associated to $p_{k}$ and $\tilde{L}^{n-1}=\tilde{L}_{1}\boxtimes\cdots\boxtimes\tilde{L}_{n-1}$. Then putting $\tilde{E}=L^{(m^{2}-m)/2}\boxtimes\tilde{L}^{n-1}$ we see that its unit sphere bundle $S(\tilde{E})$ is isomorphic to $\tilde{q}$ and therefore by combining with the equation of (8) we have
(9) $(S(\tilde{E}),\pi,(S^{2})^{(m^{2}-m)/2}\times T^{n-1})\cong\psi^{*}(S(E),\pi,G/S).$
Putting $(T_{k})^{\circ}=T_{k}-\{1\}\times S^{1}$ we write $(T^{n-1})^{\circ}=(T_{1})^{\circ}\times\cdots\times(T_{n-1})^{\circ}$. Then
@@
The restriction of $\psi$ to $((S^{2})^{(m^{2}-m)/2})^{\circ}\times(T^{n-1})^{\circ}$ is injective.
@@
In terms of the notation of the proof of Lemma 1 we prove that if we suppose that $\psi(x,y)=\psi(x^{\prime},y^{\prime})$, then it follows that $(r_{i;j},w_{i;j})_{R}=(r^{\prime}_{i;j},w^{\prime}_{i;j})_{R}$ $(0\leq j\leq m-2)$ and $(z_{2k-1},\zeta_{2k})=(z^{\prime}_{2k-1},\zeta^{\prime}_{2k})$ $(1\leq k\leq n-1)$, the latter of which is the part added to the proof of injectivity of the restriction map of $\phi$. The proof can be proceeded along the same lines as in the proof of Lemma 1 based on the result there. But in fact there are the following changes in the use of the elements $\bar{w}_{m_{j};j}$ each of which is a starting point of the proof for $\phi_{j}$: $\bar{w}_{m-1;0}\to\bar{w}_{m-1;0}$ (i.e. no change), $\bar{w}_{m-2:1}\to z_{1}\bar{w}_{m-2:1}$, $\bar{w}_{m-3:2}\to(\bar{z}_{1}\zeta_{2})\bar{w}_{m-3:2}$ and in general $\bar{w}_{m-2k-2;2k+1}\to(z_{2k+1}\bar{\zeta}_{2k})\bar{w}_{m-2k-2;2k+1}$, $\bar{w}_{m-2k-3;2k+2}\to(\bar{z}_{2k+1}\zeta_{2k+2})\bar{w}_{m-2k-3;2k+2}$ for $k=1,\ldots,n-2$. Similarly substituting these elements into the equations $\underline{a}_{0;j}=\underline{a^{\prime}}_{0;j}$, $\underline{a}_{1;j}=\underline{a^{\prime}}_{1;j}$, $\ldots$, $\underline{a}_{m-j-1;j}=\underline{a^{\prime}}_{m-j-1;j}$ we can obtain $(x,y)=(x^{\prime},y^{\prime})$.
First from $\bar{w}_{m-1;0}=\bar{w}^{\prime}_{m-1;0}$ we have $x_{0}=x^{\prime}_{0}$, which shows that $z_{1}\bar{w}_{m-2:1}=z^{\prime}_{1}\bar{w}^{\prime}_{m-2:1}$, so we have $z_{1}r_{m-2:1}=z^{\prime}_{1}r^{\prime}_{m-2:1}$. Since $r_{m-2:1}>0$, from the second equation it follows that $r_{m-2:1}=r^{\prime}_{m-2:1}$ and so $z_{1}=z^{\prime}_{1}$. This concludes that $x_{1}=x^{\prime}_{1}$. $z_{1}=z^{\prime}_{1}$, Next from $(\bar{z}_{1}\zeta_{2})\bar{w}_{m-3:2}=(\bar{z}^{\prime}_{1}\zeta^{\prime}_{2})\bar{w}^{\prime}_{m-3:2}$ we have $\zeta_{2}\bar{w}_{m-3:2}=\zeta^{\prime}_{2}\bar{w}^{\prime}_{m-3:2}$ and so it follows that $x_{2}=x^{\prime}_{2}$, $\zeta_{2}=\zeta^{\prime}_{2}$. Therefore by $(z_{3}\bar{\zeta}_{2})\bar{w}_{m-4;3}=(z^{\prime}_{3}\bar{\zeta}^{\prime}_{2})\bar{w}^{\prime}_{m-4;3}$ we have $z_{3}\bar{w}_{m-4;3}=z^{\prime}_{3}\bar{w}^{\prime}_{m-4;3}$ and so in a similar way to the previous case we have $x_{3}=x^{\prime}_{3}$, $z_{3}=z^{\prime}_{3}$. Subsequently also by $(\bar{z}_{3}\zeta_{4})\bar{w}_{m-5;4}=(\bar{z}^{\prime}_{3}\zeta^{\prime}_{4})\bar{w}^{\prime}_{m-5;4}$ we have $\zeta_{4}\bar{w}_{m-5;4}=\zeta^{\prime}_{4}\bar{w}^{\prime}_{m-5;4}$ and so it follows that $x_{4}=x^{\prime}_{4}$, $\zeta_{4}=\zeta^{\prime}_{4}$. At this point we can conclude that $x_{1}=x^{\prime}_{1}$, $x_{2}=x^{\prime}_{2}$, $x_{3}=x^{\prime}_{3}$, $x_{4}=x^{\prime}_{4}$ and $y_{1}=y^{\prime}_{1}$, $y_{2}=y^{\prime}_{2}$. These tell us that through repetition of these procedures we can arrive at the desired result. ∎
@@
Let $m=2n$ and put $M=(S^{2})^{(m^{2}-m)/2}\times T^{n-1}$. Then obviously $\dim M=\dim G/S=4n^{2}-2$. By construction we know  [4] that $\psi$ can be deformed to an onto map. Taking into account the injectivity result given in Lemma 3 together with this fact we can conclude that $\psi$ is a degree one map, that is, $\psi_{*}([M])=[G/S]$ where [][\ ][ ] denotes the fundamental class. So we have
$\langle c_{1}(E)^{2n^{2}-1},\,[G/S]\rangle=\langle(c_{1}(\psi_{*}E)^{2n^{2}-1}),\,[M]\rangle.$
Hence by (9)
$\begin{split}\langle c_{1}(E)^{2n^{2}-1},\,[G/S]\rangle&=\langle c_{1}(L^{2n^{2}-n}\hat{\otimes}\tilde{L}^{n-1}),\,[M]\rangle\\ &=\langle c_{1}(L^{2n^{2}-n}),\,[(S^{2})^{(m^{2}-m)/2}]\rangle\cdot\langle c_{1}(\tilde{L}^{n-1}),\,[T^{n-1}]\rangle\end{split}$
Substituting this into the equation of Proposition 2.1 of  [6] we obtain